OpenStudy (anonymous):

indefinite integral of x(ln x)^3 dx

OpenStudy (anonymous):

\[\int\limits x \left( \ln x \right)^3 dx=\int\limits \left( \ln x \right)^3 x~dx\] integrate by parts.

OpenStudy (anonymous):

repeat again and again to come to the function of free of ln x.